# A Survey on Code Generation with LLM-based Agents

### v1

@@ 3.2 Large Language Models
In the field of code generation, since training data contains a large amount of high-quality open-source code libraries and programming documentation, LLMs can learn and master syntactic rules of multiple programming languages as well as common programming paradigms, while understanding the mapping relationship between natural language descriptions and code logic, thereby achieving generation of executable code from natural language descriptions, or intelligent completion and refactoring in existing contexts. Typical code generation LLMs include Codex [51], CodeLlama [52], DeepSeek-Coder [53], and Qwen2.5-Coder [54], etc., which have been widely applied in software engineering scenarios such as code completion, test code generation, and bug fixing, demonstrating powerful code generation and understanding capabilities.
In addition to basic language modeling capabilities, LLMs also exhibit various important emergent abilities. First is planning capability, where LLMs can generate natural language plans to guide subsequent generation [55, 56]. Second is tool usage capability, where LLMs can actively invoke external tools based on task requirements to enhance problem-solving capabilities. By encapsulating available tool API calls, existing work has been able to integrate various external tools, such as search engines [57], calculators [58], and compilers [59], etc., to further expand the capability boundaries of LLMs. Third is environmental interaction capability, where LLMs have the ability to receive feedback from external environments and execute operations according to instructions, enabling perception, decision-making, and action in dynamic environments. The emergence of these advanced capabilities has led to the creation of agents, enabling LLMs to evolve from simple text generation tools to agent systems capable of completing more open and complex tasks.
@@ 4.1.1 Planning and Reasoning Techniques
Building on the above methods, PlanSearch [56] first formalized the planning process as an explicit search task, generating multiple sets of candidate plans and conducting parallel evaluation during the reasoning phase to find optimal solutions within a larger solution space. In terms of planning structure, both CodeTree [81] and Tree-of-Code [82] extended previous linear structures to tree structures. CodeTree organizes the code generation process into a unified tree structure that explicitly models the sequential stages of strategy exploration, solution generation, and iterative refinement. By incorporating execution-based feedback at each stage, it enables dynamic scoring and heuristic-guided pruning to efficiently navigate the large search space. Tree-of-Code explores multiple potential paths through breadth-first generation strategies and combines execution signals for branch pruning. Subsequently, multi-stage guided coding methods emphasize dividing the planning process into multi-stage control, introducing hierarchical objectives and intermediate reward signals to alleviate goal drift problems in end-to-end generation processes. Based on this, DARS [83] further adopts adaptive tree structures to enhance the decision process, branching new planning paths at key decision nodes of original planning paths based on code execution feedback, and dynamically selecting better planning paths by combining historical trajectories with execution results. Additionally, for hardware tasks, VerilogCoder [84] introduced graph-structured planning mechanisms and waveform tracing tools based on abstract syntax trees, supporting structural modeling and semantic verification of Verilog code, demonstrating the adaptive potential of planning paradigms in cross-modal and domain-specific tasks. To address the challenge that traditional search methods are difficult to apply in non-serializable environments, Guided Search [85] proposes two complementary strategies, namely one-step lookahead and trajectory selection. Both strategies are guided by a learned action-value estimator, where the model predicts potential next-step outcomes to evaluate candidate actions, and selects promising solution trajectories based on their empirical success rates. This enables effective exploration without relying on environment state serialization.
@@ 5 Applications of LLM-based Code Generation Agents in Software Development Tasks
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=center, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=center, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=7.0em,font=,, where level=2text width=7.0em,font=,, where level=3text width=7.0em,font=,, where level=4text width=7.0em,font=,, [ Applications of LLM-based
Code Generation Agents [ Automated Code
Generation and
Implementation 5.1 [ Function-Level
Code Generation [ Self-Planning [10], LATS [115], Lemur [117], CodeChain [11], MapCoder [102],
FlowGen [100], PairCoder [99], CodeTree [81], CodeCoR [12],
QualityFlow [26], CodeSim [118], DARS [83], SEW [106] , leaf, text width=25em ] ] [ Repository-Level
Code Generation [ Self-Collabration[15], ChatDev[16], Webagent[78], MetaGPT[17],
CodePlan[79], CodeAgent[28], CodePoRi[13], GameGPT[110], CodeS[119],
SoA[27], ToolGen[29], AgileCoder[120], AgileAgent[14], , leaf, text width=25em ] ] ] [ Automated
Debugging and
Program Repair 5.2 [ Self-Refine[74], Self-Debug[94], Self-Edit[75], Self-Repair[95], RepairAgent[121],
AutoCodeRover[122], SWE-Agent[123], MAGIS[98], AutoSafeCoder[124], SWE-Search[125],
HyperAgent[71], SQLFixAgent[126], OrcaLoca[127], PatchPilot[128], Thinking-Longer[129],
AdverIntentAgent[130], Nemotron-CORTEXA[131] , leaf, text width=30em ] ] [ Automated Test
Code Generation5.3 [ Automated Test
Case Generation [ TestPilot[31], CANDOR[32], XUAT-Copilot[20], LogiAgent[21],
SeedMind[35], ACH[36] , leaf, text width=25em ] ] [ Automated
Execution and
Analysis [ AUITestAgent[132], HEPH[133] , leaf, text width=25em ] ] ] [ Automated Code
Refactoring and
Optimization 5.4 [ Structural Code
Refactoring [ DataClump-Pipeline[22], iSMELL[134], EM-assist[135], HaskellAgent[136] , leaf, text width=25em ] ] [ Code Performance
Optimization [ AIDE[137], MARCO[105], LASSI-EE[33], SysLLMatic[34] , leaf, text width=25em ] ] ] [ Automated
Requirement
Clarification 5.5 [ MARE[116], ClarifyGPT[19], TiCoder[138], SpecFix[139], InterAgent[140], HiLDe[141] , leaf, text width=30em ] ] ]
@@ 5.3 Automated Test Code Generation
In unit testing, traditional search methods like EvoSuite typically rely on structural information and code execution paths, unable to understand function semantics, and can only attempt to collide with boundaries through extensive coverage or heuristic exploration. In contrast, LLMs can better understand functional intentions and semantic boundaries. TestPilot [31] can automatically generate test cases for JavaScript APIs, achieving 52.8% branch coverage on 25 npm packages and 1,684 API functions, improving 27% compared to the feedback-driven tool Nessie. CANDOR [32] can generate Java unit test cases end-to-end, successfully surpassing EvoSuite in metrics such as line coverage. In integration testing, XUAT-Copilot [20] can automatically complete the determination of test instruction types, specific parameter filling, and effect verification. The system achieves automated integration testing for mobile payment applications such as WeChat Pay. LogiAgent [21] integrates components for test scenario generation, API request execution and verification, and logic notification. It overcomes the problem that traditional REST API testing only focuses on detecting server crashes and error codes. Meanwhile, this system achieves automated detection of a series of logic problems brought by business development and domain-specific requirements. In security testing, LLM-based code generation agents can automatically derive aggressive input patterns and boundary condition cases from natural language documentation, interface definitions, and code structures. SeedMind [35] automatically generates high-quality seed inputs through LLMs as starting points for gray-box fuzzing. The framework can dynamically adjust generation strategies based on test feedback to improve test coverage. Meta’s ACH [36] system generates mutants that conform to program semantics through LLMs and judges mutant equivalence, improving the efficiency of mutation testing.
Automated testing processes involve not only test case generation but also execution, evaluation, and analysis. For example, AUITestAgent [132] can automatically identify interactive instructions and execute full-process graphical user interface operations and functional verification through pure natural language testing requirements. The HEPH system from NVIDIA achieves verification of correctness of C/C++-based embedded systems in various scenarios by analyzing requirement documents and retrieving various related files, and guides the next round of test case generation with coverage reports after each testing round.
@@ 5.4 Automated Code Refactoring and Optimization
In code performance optimization, both LASSI-EE [33] and SysLLMatic [34] use LLMs to analyze various performance and energy consumption metrics during code runtime, achieving improvements in software system performance and energy consumption through multi-round diagnosis and optimization.
@@ 6.2 Evaluation Metrics
[…]
@@ 7 Deployed Code Generation Agent Tools
Several multi-agent code generation tools have been deployed and have produced widespread impact in the current market. We roughly categorize them into three types to show their evolutionary trajectory: 1) Co-pilot: Human-machine close collaboration, serving as the co-pilot of developers to assist programmers in coding. 2) Collaborator: Capable of understanding entire codebase contexts and engaging in deep interactive collaboration with developers. 3) Autonomous Team: Aimed at automating the entire development process, where humans play more of a client or manager role. Currently, emerging code generation agent tools in the market embody this evolutionary trajectory, playing roles at different stages of programming assistant, collaborative partner, and autonomous development team.
@@ 8 Challenges and Future Directions
We discuss and summarize the challenges currently faced by LLM-based code generation agents and future development directions from four dimensions.
@@ 8.1 Limitations of Agent Core Capabilities
Context Engineering: In practical scenarios, agents often face scattered and incomplete information across multiple files, modules, or documentation. Without well-organized context, they may generate inconsistent or incorrect code. This problem becomes more severe in multi-step tasks, where key information from earlier steps can be easily lost or misunderstood. While some recent approaches try to improve this through retrieval or memory mechanisms, there is still no robust solution that works reliably across large and complex codebases.
@@ 8.2 Robustness and Updatability Challenges of Agent Systems
Agent Knowledge Updates and Continuous Learning: Software development environments are dynamically evolving, but agents constructed through one-time training have fixed knowledge bases that inevitably become outdated. Agents need capabilities to actively and continuously learn unique coding standards and domain knowledge from specific projects. However, there is currently a lack of effective continuous learning mechanisms that can efficiently integrate new knowledge into agents while avoiding catastrophic forgetting.
8.3 Human-Agent Interaction, Trustworthiness, and Cost Challenges
Reliability and Debuggability of Agent Systems: The inherent hallucination problem of LLMs easily causes them to generate seemingly credible but actually false hallucination content, directly undermining the reliability of agent systems. If agent systems cannot consistently and reliably provide accurate information, they will be extremely dangerous in high-risk scenarios (such as critical infrastructure control). Additionally, when agent system behavior deviates from expectations, their complex internal decision chains and non-deterministic outputs make debugging and repairing problems difficult.
Flexibility and Security in Tool Usage: To perform meaningful tasks, agents need access to external tools such as compilers, test frameworks, and database interfaces. However, relying on a predefined and static toolset limits the flexibility of agents and makes it difficult to handle diverse and evolving real-world scenarios. A key challenge is to enable agents to move beyond these limitations by supporting flexible and on-demand tool discovery and integration. Meanwhile, it is essential to ensure the security of the tool execution process to prevent unintended or malicious behavior.
@@ 8.3 Human-Agent Interaction, Trustworthiness, and Cost Challenges
High Operating Costs and Performance Bottlenecks: Multi-agent systems improve problem-solving capabilities through multi-round collaborative interactions, but this typically comes at the cost of high computational and time expenses. Each interaction between agents and each call to LLMs generates significant computational and time overhead. When tasks are complex and interaction rounds increase, total costs rise sharply, becoming a major economic and technical barrier limiting large-scale deployment of multi-agent systems in real development scenarios.

### v2

@@ 3.2 Large Language Models
In the field of code generation, since training data contains a large amount of high-quality open-source code libraries and programming documentation, LLMs can learn and master syntactic rules of multiple programming languages as well as common programming paradigms, while understanding the mapping relationship between natural language descriptions and code logic, thereby achieving generation of executable code from natural language descriptions, or intelligent completion and refactoring in existing contexts. Typical code generation LLMs include Codex [51], CodeLlama [rozière2024codellamaopenfoundation], DeepSeek-Coder [52], and Qwen2.5-Coder [53], etc., which have been widely applied in software engineering scenarios such as code completion, test code generation, and bug fixing, demonstrating powerful code generation and understanding capabilities.
In addition to basic language modeling capabilities, LLMs also exhibit various important emergent abilities. First is planning capability, where LLMs can generate natural language plans to guide subsequent generation [54, 55]. Second is tool usage capability, where LLMs can actively invoke external tools based on task requirements to enhance problem-solving capabilities. By encapsulating available tool API calls, existing work has been able to integrate various external tools, such as search engines [56], calculators [57], and compilers [58], etc., to further expand the capability boundaries of LLMs. Third is environmental interaction capability, where LLMs have the ability to receive feedback from external environments and execute operations according to instructions, enabling perception, decision-making, and action in dynamic environments [59, 60] .
@@ 4.1.1 Planning and Reasoning Techniques
Building on the above methods, PlanSearch [55] first formalized the planning process as an explicit search task, generating multiple sets of candidate plans and conducting parallel evaluation during the reasoning phase to find optimal solutions within a larger solution space. In terms of planning structure, both CodeTree [82] and Tree-of-Code [83] extended previous linear structures to tree structures. CodeTree organizes the code generation process into a unified tree structure that explicitly models the sequential stages of strategy exploration, solution generation, and iterative refinement. By incorporating execution-based feedback at each stage, it enables dynamic scoring and heuristic-guided pruning to efficiently navigate the large search space. Tree-of-Code explores multiple potential paths through breadth-first generation strategies and combines execution signals for branch pruning. Subsequently, Multi-stage guided code generation for Large Language Models [84] emphasize dividing the planning process into multi-stage control, introducing hierarchical objectives and intermediate reward signals to alleviate goal drift problems in end-to-end generation processes. Based on this, DARS [85] further adopts adaptive tree structures to enhance the decision process, branching new planning paths at key decision nodes of original planning paths based on code execution feedback, and dynamically selecting better planning paths by combining historical trajectories with execution results. Additionally, for hardware tasks, VerilogCoder [86] introduced graph-structured planning mechanisms and waveform tracing tools based on abstract syntax trees, supporting structural modeling and semantic verification of Verilog code, demonstrating the adaptive potential of planning paradigms in cross-modal and domain-specific tasks. To address the challenge that traditional search methods are difficult to apply in non-serializable environments, Guided Search [87] proposes two complementary strategies, namely one-step lookahead and trajectory selection. Both strategies are guided by a learned action-value estimator, where the model predicts potential next-step outcomes to evaluate candidate actions, and selects promising solution trajectories based on their empirical success rates. This enables effective exploration without relying on environment state serialization.
@@ 5 Applications of LLM-based Code Generation Agents in Software Development Tasks
[…]
Applications of LLM-based
Code Generation Agents
Automated Code Generation and Implementation 5.1 Function-Level Code Generation Self-Planning [10], LATS [117], Lemur [119], CodeChain [11], MapCoder [104], FlowGen [102], PairCoder [101], CodeTree [82], CodeCoR [12], QualityFlow [26], CodeSim [120], DARS [85], SEW [108] Repository-Level Code Generation Self-Collabration[15], ChatDev[16], Webagent[79], MetaGPT[17], CodePlan[80], CodeAgent[28], CodePoRi[13], GameGPT[112], CodeS[121], SoA[27], ToolGen[29], AgileCoder[122], AgileAgent[14], Automated Debugging and Program Repair 5.2 Self-Refine[75], Self-Debug[96], Self-Edit[76], Self-Repair[97], RepairAgent[123], AutoCodeRover[124], SWE-Agent[125], MAGIS[100], AutoSafeCoder[126], SWE-Search[127], HyperAgent[72], SQLFixAgent[128], OrcaLoca[129], PatchPilot[130], Thinking-Longer[131], AdverIntentAgent[132], Nemotron-CORTEXA[133] Automated Test Code Generation5.3 Automated Test Case Generation TestPilot[31], CANDOR[32], XUAT-Copilot[20], LogiAgent[21], SeedMind[35], ACH[36] Automated Execution and Analysis AUITestAgent[134], HEPH[135] Automated Code Refactoring and Optimization 5.4 Structural Code Refactoring DataClump-Pipeline[22], iSMELL[136], EM-assist[137], HaskellAgent[138] Code Performance Optimization AIDE[139], MARCO[107], LASSI-EE[33], SysLLMatic[34] Automated Requirement Clarification 5.5 MARE[118], ClarifyGPT[19], TiCoder[140], SpecFix[141], InterAgent[142], HiLDe[143]
@@ 5.3 Automated Test Code Generation
In unit testing, traditional search methods like EvoSuite [145] typically rely on structural information and code execution paths, unable to understand function semantics, and can only attempt to collide with boundaries through extensive coverage or heuristic exploration. In contrast, LLMs can better understand functional intentions and semantic boundaries. TestPilot [31] can automatically generate test cases for JavaScript APIs, achieving 52.8% branch coverage on 25 npm packages and 1,684 API functions, improving 27% compared to the feedback-driven tool Nessie. CANDOR [32] can generate Java unit test cases end-to-end, successfully surpassing EvoSuite in metrics such as line coverage. In integration testing, XUAT-Copilot [20] can automatically complete the determination of test instruction types, specific parameter filling, and effect verification. The system achieves automated integration testing for mobile payment applications such as WeChat Pay. LogiAgent [21] integrates components for test scenario generation, API request execution and verification, and logic notification. It overcomes the problem that traditional REST API testing only focuses on detecting server crashes and error codes. Meanwhile, this system achieves automated detection of a series of logic problems brought by business development and domain-specific requirements. In security testing, LLM-based code generation agents can automatically derive aggressive input patterns and boundary condition cases from natural language documentation, interface definitions, and code structures. SeedMind [35] automatically generates high-quality seed inputs through LLMs as starting points for gray-box fuzzing. The framework can dynamically adjust generation strategies based on test feedback to improve test coverage. Meta’s ACH [36] system generates mutants that conform to program semantics through LLMs and judges mutant equivalence, improving the efficiency of mutation testing.
Automated testing processes involve not only test case generation but also execution, evaluation, and analysis. For example, AUITestAgent [134] can automatically identify interactive instructions and execute full-process graphical user interface operations and functional verification through pure natural language testing requirements. The HEPH system [135] from NVIDIA achieves verification of correctness of C/C++-based embedded systems in various scenarios by analyzing requirement documents and retrieving various related files, and guides the next round of test case generation with coverage reports after each testing round.
@@ 5.4 Automated Code Refactoring and Optimization
In code performance optimization, both LASSI-$E$ [33] and SysLLMatic [34] use LLMs to analyze various performance and energy consumption metrics during code runtime, achieving improvements in software system performance and energy consumption through multi-round diagnosis and optimization.
@@ 6.2 Evaluation Metrics
$\text{pass@k}=\mathbb{E}_{\text{problems}}\left[1-\frac{\binom{n-c}{k}}{\binom{n}{k}}\right$]
@@ 7 Deployed Code Generation Agent Tools
Several multi-agent code generation tools have been deployed and have produced widespread impact in the current market. We roughly categorize them into three types to show their evolutionary trajectory: 1) Co-pilot: Human-machine close collaboration, serving as the co-pilot of developers to assist programmers in coding. 2) Collaborator: Capable of understanding entire codebase contexts and engaging in deep interactive collaboration with developers. 3) Autonomous Team: Aimed at automating the entire development process, where humans play more of a client or manager role. Currently, emerging code generation agent tools in the market embody this evolutionary trajectory, playing roles at different stages of programming assistant, collaborative partner, and autonomous development team. Table I presents a horizontal comparison of these mainstream tools.
TABLE I: Horizontal Comparison of Code Generation Agent Tools
Tool Name Core Paradigm Primary Integration Method Context Engine Best Application Scenario Main Advantages Known Limitations
GitHub Copilot Programming Assistant IDE Extension Plugin Retrieval Augmented Code Completion Wide adoption, excels at routine tasks Requires human supervision, may overlook global context
Devin Autonomous Development Team Independent Online Platform Sandbox-based Knowledge Base End-to-End Code Generation High autonomy, handles full task lifecycle Low practical reliability, unpredictable results, prone to loops
Cursor Collaborative Partner AI-Native IDE Vector Semantic Indexing Code Completion, Code Development, Codebase Q&A Deep codebase understanding, excellent interaction experience Weaker autonomous agent capabilities
Tongyi Lingma Collaborative Partner Enterprise Platform Integration Repository Knowledge Graph Code Development Robust and systematic exploration of solution space More focused on enterprise-level applications
Claude Code Autonomous Development Team Command Line Interface Ultra-Long Context Window + Agent Search End-to-End Code Generation, Codebase Q&A Powerful code agent, global codebase understanding Potentially high cost
@@ 8 Challenges and Future Directions
We discuss and summarize the challenges currently faced by LLM-based code generation agents and future development directions from five dimensions.
@@ 8.1 Limitations of Agent Core Capabilities
Context Engineering: In real-world software development scenarios, information is often scattered across multiple files, modules, and documents. Without a systematic mechanism for constructing and managing context, models are highly prone to generating inconsistent or even incorrect code. This issue becomes particularly prominent in multi-step tasks. Empirical evidence suggests that many failures of intelligent agents do not stem from the models themselves, but rather from defects in the ingested context. These defects can be categorized into four typical problems: context poisoning (incorrect or hallucinated information contaminates subsequent reasoning), context distraction (redundant information overwhelms key signals), context confusion (format inconsistencies lead to misinterpretation), and context conflict (contradictory information causes decision-making errors). Although existing studies have attempted to mitigate this challenge through approaches such as retrieval augmentation and long-term memory modules, there is still a lack of stable and practical solutions for software development. Therefore, context engineering is not merely an extension of prompt design, but rather a dynamic methodology whose core objective is to deliver the right information and tools to the model—at the right time and in the appropriate format. Only through systematic organization and optimization of context can intelligent agents maintain consistency, reliability, and efficiency when operating in complex tasks and large-scale systems.
@@ 8.2 Robustness and Updatability Challenges of Agent Systems
Agent Knowledge Updates and Continuous Learning: Software development environments are inherently dynamic, with new programming language features and framework versions constantly emerging. However, code-generation agents trained in a one-off manner possess fixed knowledge bases, making them unable to keep up with the latest technologies and inevitably prone to obsolescence. Thus, agents need the capability to actively and continually acquire new knowledge. Efficient fine-tuning methods such as LoRA [167] can incrementally adjust model parameters, but their effectiveness in learning new knowledge remains limited and they inevitably suffer from degradation of previously acquired knowledge. Retrieval-augmented generation (RAG), on the other hand, relies on external vector databases for real-time retrieval, yet it requires continuous maintenance of external knowledge stores and lacks deep integration between retrieved content and the internal knowledge of the model. Therefore, an effective continual learning mechanism that can efficiently incorporate new knowledge into agents while avoiding catastrophic forgetting is still missing.
8.3 Challenges of Tool Integration and Deployment for Agents in Open Environments
Flexibility and Security in Tool Usage: To accomplish meaningful tasks, agents must be equipped with the ability to use external tools (e.g., compilers, testing frameworks, database interfaces). However, providing agents with a predefined and static toolset lacks flexibility and cannot meet the needs of open and dynamic real-world tasks. Thus, an important research question is how to enable agents to move beyond the constraints of static tools, achieving on-demand and flexible tool discovery and integration, while simultaneously ensuring the security of tool execution to prevent potential malicious operations.
@@ 8.3 Challenges of Tool Integration and Deployment for Agents in Open Environments
High Operational Costs of Agent Systems: Multi-agent systems enhance problem-solving capabilities through iterative collaboration and interaction, but this often comes at the expense of significant computational and time costs. Each interaction among agents and every call to a large language model incurs substantial overhead. As task complexity increases and interaction rounds grow, the overall cost escalates sharply, becoming a major economic and technical barrier to large-scale deployment of multi-agent systems in real-world development scenarios.
Lightweight Deployment on Edge Devices: Deploying code-generation agents on edge devices such as smartphones and drones can reduce reliance on cloud servers, better satisfying needs of users for real-time response and autonomy. However, the limited computational and storage capacities of such devices pose challenges for lightweight deployment of agent systems. Future solutions will require model compression and inference acceleration techniques to reduce computational and storage overhead while preserving generation capabilities. Examples include quantization to lower representation costs of model weights, knowledge distillation to transfer capabilities from large models to smaller ones, and inference pruning to shorten inference length.
8.4 Trustworthiness, Security, and Ethical Risks of Agent Systems
Reliability and Debuggability of Agent Systems: The inherent hallucination problem of large language models makes them prone to generating content that appears plausible but is in fact false, thereby undermining the reliability of agent systems. If agent systems cannot consistently and reliably provide accurate information, they become extremely dangerous in high-risk scenarios (e.g., critical infrastructure control). Furthermore, when agent behavior deviates from expectations, their complex internal decision-making chains and nondeterministic outputs make debugging and troubleshooting particularly difficult.
Malicious Code Generation: As the capabilities of agent systems improve, their security risks require greater attention. On the one hand, users may craft prompts to bypass model safety mechanisms and induce the generation of malicious code. On the other hand, agents may memorize and reproduce historical data containing security vulnerabilities or attack scripts, thereby automatically generating malicious code targeting specific systems. To mitigate such risks, comprehensive security review mechanisms must be introduced before and after the generation process [168]. Potential measures include conducting interpretability analyses of generation paths to identify risks prior to code output, and applying sandbox-based detection to sensitive operations in the generated content afterward.
Code Copyright and Ownership: Code-generation agents are trained on public code data from open-source repositories, and may also involve private code data from enterprises and individuals. Consequently, agents may generate code segments that closely resemble copyright-protected code, or reproduce open-source content subject to licensing restrictions that is then used for commercial purposes—leading to copyright infringement risks. Moreover, due to the current lack of unified standards for copyright recognition of agent-generated content, the legal responsibilities associated with using such systems in software development remain ambiguous and prone to disputes. Future research needs to incorporate techniques such as “digital watermarking” [169] during both training and inference to trace the origin of generated content, alongside establishing robust community-level ethical guidelines.